# ECGrecover: a Deep Learning Approach for Electrocardiogram Signal Completion

### v1

@@ 3. Results
In this section, we evaluate the effectiveness of the proposed method, ECGrecover. We first outline the experimental setup and then we discuss the results. Additional simulations and findings are provided in Appendix A for space reasons.
@@ 3.1.1. Datasets
We utilized the Generepol dataset (Salem et al., 2017), comprising 10-second 8-lead ECGs sampled at 500Hz. Leads III, aVF, aVL, and aVR were computed from the eight other leads I, II, and V1-V6, following the common practice in the ECG literature (Salem et al., 2017),
The study was conducted at the Clinical Investigation Center of the Pitié-Salpétrière hospital from 2008 to 2012 in Paris, France and involved 15 119 ECG recordings from 990 healthy subjects. These recordings were taken before and at 1, 2, 3, and 4 hours after the administration of 80mg oral sotalol dose, denoted as sotT1, sotT2, sotT3, and sotT4, respectively. To ensure data quality, two expert cardiologists curated the recordings. We refer to (Salem et al., 2017; Prifti et al., 2021) for further details about the acquisition process.
The raw dataset used for training denoted as in Section 2.1, comprises 4 498 ECG recordings. When augmented through the application of various masks as detailed in Figure 2, this dataset expands to a total of 76 466 samples. For the validation set, the original 558 ECG recordings are augmented to 9 486 ECG. Similarly, the testing set, initially consisting of 591 ECGs, is augmented to a total of 10 047 ECG
@@ 3.1.2. Data preprocessing
We first min-max normalized the ECG data to a range of [-1,1]. For the preparation of our data, we followed a strategy close to (Joo et al., 2023). In short, we applied bandpass filtering to the normalized data, using lower and upper cut-off frequencies of 0.05Hz and 150Hz, respectively, to eliminate noise such as baseline drift and improve signal quality. Finally, we down-sampled the original signal from 5000pts (500Hz) to 512pts (50Hz)
As described in Section 2.1, to mimic real-life cases of ECG with incomplete leads or segments, we considered the scenarii illustrated in Figure 2. For a given 12-lead ECG input signal, we initially segmented the signal following the different configurations at coordinates corresponding to the primers. Subsequently, to maintain the original dimensions of the signal, we completed the primers with random noise. This noise is generated from a continuous uniform distribution within the interval. This method introduces variability in the noise across samples, thereby realistically simulating the presence of missing signals. An illustration of how an ECG with incomplete leads or segments would appear after introducing random noise is shown in Figure 3 top.
@@ 3.1.3. Evaluation metrics
We assessed the effectiveness of the proposed approach from two distinct perspectives. Initially, we evaluated its reconstruction capability using standard distortion metrics employed in recent literature (Joo et al., 2023; Seo et al., 2022). These metrics measure the reconstruction error between the original and reconstructed signals, with lower values indicating better performance. Specifically, we employed the root mean square error (RMSE) and the maximum absolute error (MAE). Additionally, we utilized the dynamic time warping (DTW) as an alignment-based similarity measure to evaluate temporal alignment between the original and reconstructed signals, favoring lower values. The Pearson correlation coefficient (PCC) evaluates linear relationships between leads, with higher values indicating better alignment.
Subsequently, we examined ECGrecover ability to preserve crucial features of the ECG signal. In particular, we focused on preserving wave coordinates, including peaks, i.e., P, Q, R, S, T, as illustrated in Figure 9. We used the neurokit2 python library for this task (Makowski et al., 2021). The results were evaluated by measuring the difference in seconds between the positions of the peaks in the original ECG and the reconstructed one, denoted as (the lower, the better), where represents one of the potential peaks of interest.
@@ 3.1.4. State-of-the-art methods
We compared ECGrecover with the most recent state-of-the-art (SOTA) methods for the ECG lead-reconstruction problem333No comparison could be made with Seo et al. (2022) (Seo et al., 2022) due to the unavailability of their code.. It’s worth noting that, to the best of our knowledge, no existing methods address the ECG segment-recovery problem. We briefly introduce these SOTA as follows:
Pix2Pix (Isola et al., 2017). A conditional generative adversarial network (cGAN), initially proposed for images, employs a 2D U-Net generator and a PatchGAN discriminator. Unlike traditional discriminators that classify the entire image as real or fake, PatchGAN evaluates the realism of small, overlapping patches within the image.
EKGAN (Joo et al., 2023). A generative adversarial network (GAN) for the reconstruction, which is based on Pix2Pix augmented by an additional label generator and the use of a 1D U-Net discriminator instead of a PatchGAN discriminator.
CopyPaste. A naive-approach, which consists in replicating the primers throughout the signal to complete the ECG. This technique is notably implemented in Badilini et al. (2005), a software used to digitize paper ECGs.
In their original paper, particularly in EKGAN, the focus is on the reconstruction of the 11 leads given lead I, corresponding to setting C in Figure 2. Hence, the method was not initially designed for the ECG segment-recovery problem. Nevertheless, for completeness and fairness, we assessed all the methods performance across all the settings in Figure 2.
@@ 3.1.5. Model architecture and training procedure
We consider a U-Net like architecture (Ronneberger et al., 2015), which is summarized in Figure 3. The encoder consists of two input branches one 2D and the other one 1D. On the one hand, the purpose of the 2D branch is to let the model learn inter-lead information of the signal. This branch is made of 4, 2D convolution blocks, each of them presenting a 2D convolution layer, batch normalization, a LeakyReLu () activation function, and dropout (). On the other hand, the 1D branch provides the model to learn intra-lead characteristics. Similarly to the 2D branch, this one consists of , 1D convolution blocks. The output of each block (1D/2D) is concatenated and then given to the decoder during the decoding phase.
The transition block consists of a 2D deconvolution layer, kernels of dimension with no stride, followed by batch normalization and a LeakyReLu () activation function. The output of the transition is fed into the decoder, which consists of 4, 2D deconvolution blocks. Each of them is composed of a 2D deconvolution layer, batch normalization, and LeakyReLu activation function (). Each output of a deconvolution block is concatenated with the output (concatenated 1D/2D) of the symmetric convolution block. Finally, in the last 2D deconvolution block of the decoder, we use the Tanh activation function to have a signal between and as in the original normalized signal in the input.
We trained the model for 100 epochs with a batch size of 256. We also used the Adam optimizer with a learning rate of 0.01. The hyper-parameter in Equation 1 was tuned by choosing values within the range . The final value of used for the experiments was fixed to (cf. Section A.1).
@@ 3.2. Discussion
We now present the main experimental results to show the effectiveness of ECGrecover for both the ECG lead-reconstruction and the ECG segment-recovery problems. In particular, we first discuss the capability of our approach by considering the metrics in Section 3.1.3; then, we analyze the ability of the method to preserve important features of the ECG; finally, we assess how the reconstructed signal affects the inter-lead correlation. Further results are provided in Appendices A and B.
Specifically, throughout this section, we use the C notation to indicate the input sample that, at testing time, has been masked according to the considered configuration . Therefore, we compare the reconstructed lead (for example, lead I from the ECG in output when the input sample was masked with the C1 configuration) with its original counterpart.
@@
Lead | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | ECGrecover | EKGAN (Joo et al., 2023) | Pix2Pix (Isola et al., 2017) | CopyPaste | |
@@
Figure 5 presents an evaluation of our proposed method, ECGrecover, alongside benchmark state-of-the-art approaches, focusing on PCC, RMSE and DTW for the Generepol dataset. We relegate to Appendix B the complete sets of the results across all the 17 different scenarii in Figure 2 and all the 12 leads.
@@
As for the ECG segment-recovery problem, in Figure 5, we specifically focused on the configuration C1 as it represents the scenario with primers of shorter duration (i.e., 0.8sec) where most of the signal needs to be reconstructed. Additionally, we analyze the settings C3 and C5 as they reflect real-life situations when digitizing paper ECGs. Regarding the ECG lead-reconstruction problem, we highlight in the table the configurations C as it is the one considered also in (Joo et al., 2023), and C because of its relevance in wearable device recordings and paper-stored ECGs, where Lead II, akin to Lead I, is predominantly captured. For each of the input masks, we tested the ability of the model to reconstruct the limb leads I, II, and III and the chest leads V1 and V6. Practitioners note that leads II and V1 contain the most information, while lead III contains the least.
@@
Overall, ECGrecover consistently outperforms the SOTA, regardless of the chosen metric. As depicted in Figure 5, lead V6 emerges as the most straightforward to reconstruct, both in ECG segment-recovery (C1, C3, C5) and ECG lead-reconstruction (CI, CII). Conversely, lead III proves to be the most challenging to reconstruct from the other leads. This can be explained by the perpendicular information that it captures relative to the other leads. In particular, while our method consistently delivers satisfactory results across all samples tested, competing methods exhibit greater variability (e.g., DTW values for EKGAN). Interestingly, the simple CopyPaste method yields reasonable results regarding RMSE and DTW. However, as seen from the PCC results, the method fails to preserve the linear relationships between the original and reconstructed leads, underscoring the method’s limitations as the heart rhythm may vary through time, thus losing sync. This highlights the importance of evaluating the methods from multiple perspectives. Relying solely on traditional distortion metrics, primarily measuring spatial distances between signals, may lead to overly optimistic results. Referring to the comprehensive results in Figure 10, our analysis shows that, in contrast to ECGrecover, the competing approaches EKGAN and Pix2Pix face significantly more challenges in the ECG segment-recovery task compared to the ECG lead-reconstruction task. On the other hand, the CopyPaste method struggles primarily with the lead-reconstruction problem. This outcome is not surprising as EKGAN and Pix2Pix were precisely designed for ECG lead-reconstruction, whereas CopyPaste, which essentially duplicates the signal across leads, performs poorly when the lead to reconstruct is less correlated with the primer lead in the mask.
Finally, we emphasize that in Section A.3, we conduct an additional evaluation of our proposed method using the PTB-XL dataset (Wagner et al., 2020). The results underscore that the effectiveness of ECGrecover is not confined to our Generepol dataset thereby demonstrating its applicability and robustness across different datasets.
@@ 3.2.2. On the QT distance preservation
An ECG visually represents the heart’s electrical activity during a cardiac cycle. This cycle, which encompasses a single heartbeat, involves the heart filling with blood and then contracting to pump blood to the rest of the body. The standard ECG trace includes several key components, each corresponding to specific electrical events within the heart. As shown in Figure 9, the signal typically includes identifiable points corresponding to the start, end and peaks of the underlying waves P and T and the QRS complex. The intervals and segments between these components (QT distance, PR interval, etc.) also provide important information about the cardiac cycle’s timing and electrical conduction through the heart. Understanding and evaluating these components helps diagnose a wide range of cardiac abnormalities and conditions (Parmet et al., 2003).
In Table 2, we evaluate the methods’ capacity to generate a signal that can be effectively analyzed. The performance is computed as the difference between the QT distance calculated in the reconstructed signals and the QT distance computed in the original signals (-QT). Specifically, we segment each ECG into individual heart-beats by identifying the position of the R-peak within the original leads. Then, we isolate the specific spike by taking a window of 50 points around the R peaks—20 points before and 30 points after. For the detection of the peaks, we use the Neurokit library444https://neuropsychology.github.io/NeuroKit/, which supports the extraction of critical features from the ECG, such as P, Q, R, S, and T peaks, by utilizing advanced signal processing techniques. At each heart-beat, we identified the Q wave and the end of the T wave, allowing us to calculate the mean QT interval for each ECG.
The results in Table 2 are grouped w.r.t. Sot+ and Sot- true label, originally assigned to the ECGs. Indeed, as described in Section 3.1.1, in the Generepol study, participants were administered Sotalol, which is known to prolong the QT interval and sometimes modify the morphology of the T-wave, which makes the measurement of QT more challenging (Salem et al., 2017; Prifti et al., 2021). The ECG recordings were taken before (Sot-) and after (Sot+) the drug administration.
Remarkably, our analysis did not reveal a significant performance difference of ECGrecover between the pre-sotalol and post-sotalol ECGs. The methods demonstrate consistent performance across various input masks and lead reconstructions. While ECGrecover consistently outperforms its competitors, the CopyPaste method occasionally achieves better results. The explanation lies in the inherent nature of the CopyPaste technique, which maintains beat information through direct replication. By duplicating each beat identically, the CopyPaste method inherently preserves the average QT distance, with a lower variability.
@@
We conclude the assessment of ECGrecover by analyzing the inter-lead correlation. Specifically, in Figure 6, the matrix entry at index represents PCC between the -th reconstructed lead and the -th original lead, given that the input mask applied is . Thus, the -th row illustrates the model’s effectiveness in utilizing the specific -th lead’s data to accurately reconstruct the remaining leads of the ECG. Conversely, the -th column indicates which input mask most significantly guides the model towards a more accurate reconstruction.
Figure 6 highlights the strong correlation between leads V3, V4, V5 and V6. Lead III exhibits a lower correlation with the other leads. While surprising at first, given that lead III can theoretically be derived from leads I and II through Einthoven’s triangle (Parmet et al., 2003), this characteristic has also been observed in (Jain et al., 2023), where a similar analysis was conducted on a distinct dataset. Indeed, lead III captures a perpendicular view of the heart compared to the others.
@@
Moreover, similar conclusions can also be found in (Zhang et al., 2022), regarding lead V1. Indeed, Zhang et al. demonstrated that the absence of lead V1 could significantly affect the diagnostic accuracy of various deep neural network models. Their findings revealed a notably low correlation between lead V1 and the other leads, with PCC close to 0. This low correlation suggests that the information provided by lead V1 is substantially distinct from that of the other leads. Consequently, it is understandable that the mask C in Figure 6 is the most challenging to reconstruct. Noteworthy, PCC measures linear correlation, and if there are nonlinear relationships or complex dependencies between these leads, PCC may not fully capture such associations. Factors like noise, artifacts, or variations in individual physiology can also contribute to lower correlation values.

### v3

@@ 3. Results
In this section, we evaluate the effectiveness of ECGrecover and compare it with state-of-the-art (SOTA) methods. We first outline the experimental setup, and then we discuss the results. Additional simulations and findings are provided in Appendix A and Appendix B.
@@ 3.1.1. Datasets
We utilized the Generepol dataset (salem2017genome) (cf. Table 7), comprising 10-second 8-lead ECGs sampled at 500Hz. Leads III, aVF, aVL, and aVR were computed from leads I and II, following the common practice in the ECG literature (ECGFrontales).
The study was conducted from 2008 to 2012 at the Clinical Investigation Center of the Pitié-Salpétrière Hospital Paris, France. 15119 ECGs were recorded from 989 healthy subjects. These recordings were taken before and after the 80mg of Sotalol oral intake. Two expert cardiologists curated and evaluated the recordings. We refer to (salem2017genome; prifti2021deep) for further details about the acquisition process.
The raw data for the training set consists of 4498 ECGs (i.e., in Section 2.1), augmented to 76466 after each of the masks in Figure 2 is applied. The validation set consists of 9486 ECGs, and the testing set consists of 10047 ECGs.
@@ 3.1.2. Data preprocessing
We first min-max normalized the ECG data to a range of [-1,1]. We followed a strategy close to (joo2023twelve) to prepare our data: bandpass filtering was applied to the normalized data, using lower and upper cut-off frequencies of 0.05Hz and 150Hz, respectively, to eliminate noise, such as baseline drift and improve signal quality. Finally, we downsampled the original signal from 5000pts (500Hz) to 512pts (50Hz). As described in Section 2.1, to mimic real-life cases of ECG with incomplete leads or segments, we considered the scenarios illustrated in Figure 2. Subsequently, to maintain the original dimensions of the signal, we completed the primers with random noise. Such noise was generated from a continuous uniform distribution within the interval. This method introduces variability in the noise across samples, thereby realistically simulating the presence of missing signals. An illustration of an ECG with incomplete leads or segments after introducing random noise is shown in Figure 3 (top).
@@ 3.1.3. Evaluation metrics
We assessed the effectiveness of the proposed approach from two distinct perspectives. Initially, we evaluated its reconstruction capability using standard distortion metrics employed in recent literature (joo2023twelve; seo2022multiple). We employed the root mean square error (RMSE) and the maximum absolute error (MAE). Additionally, we considered the dynamic time warping (DTW) as an alignment-based similarity measure to evaluate the distance between the original and reconstructed signals (the lower, the better) and the Pearson correlation coefficient (PCC) to evaluate the linear relationships between leads (the higher, the better). Subsequently, we examined ECGrecover’s ability to preserve crucial features of the ECG signal. In particular, we focused on preserving the QT distance. We isolated each ECG heartbeat by detecting the R peak and extracting the segment 0.4 seconds before and 0.6 seconds after. For each heartbeat, we identified the position of the Q and T peaks333We used the neurokit2 python library to detect the peaks (Makowski2021neurokit). to determine the QT distance. We then calculated the difference in QT distance between the reconstructed ECGs and the real ECGs denoted as QT (the lower, the better).
@@ 3.1.4. State-of-the-art methods
We compared ECGrecover with the most recent SOTA methods for the ECG lead-reconstruction problem444No comparison could be made with Seo et al. (2022) (seo2022multiple) due to code unavailability.. It is worth noting that, to the best of our knowledge, no other existing methods address the ECG segment-recovery problem.
Pix2Pix (isola2017image). Initially proposed for images, a conditional generative adversarial network (cGAN) employs a 2D U-Net generator and a PatchGAN discriminator. Unlike traditional discriminators that classify the entire image as real or fake, PatchGAN evaluates the realism of small, overlapping patches within the image.
EKGAN (joo2023twelve). A GAN for the reconstruction, based on Pix2Pix, augmented by an additional label generator and a 1D U-Net discriminator instead of a PatchGAN discriminator.
CopyPaste. A naive approach consists of replicating the primers throughout the signal to complete the ECG. This technique is implemented in (badilini2005ecgscan), a software for digitizing paper ECGs.
Their original paper, particularly EKGAN, focuses on reconstructing the 11 leads given lead I, corresponding to setting C in Figure 2. Hence, the method was not initially designed for the ECG segment-recovery problem. However, for fairness, we assessed all the methods’ performance across all the settings in Figure 2.
@@ 3.1.5. Model architecture and training procedure
We consider a U-Net like architecture (ronneberger2015u), summarized in Figure 3. The encoder consists of two input branches, one 2D and the other 1D. On the one hand, the purpose of the 2D branch is to let the model learn inter-lead information of the signal. The 2D branch processes all leads simultaneously, facilitating the transfer of spatial information (inter-lead). This branch comprises four 2D convolution blocks, each presenting a 2D convolution layer, batch normalization, a LeakyReLu () activation function, and dropout (). On the other hand, the 1D branch provides the model to learn intra-lead characteristics. The 1D branch focuses on lead-specific information, maintaining horizontal coherence (intra-lead). Similarly, to the 2D branch, this one consists of , 1D convolution blocks. The output of each block (1D/2D) is concatenated and then input in the decoder. The skip connection characteristics of the U-Net play an essential role in preserving part of the information captured by the 1D and 2D branches, which they transmit to the decoder. In this way, they maintain intra-lead and inter-lead relationships throughout the auto-encoding and decoding process.
The transition block consists of a 2D deconvolution layer, kernels of dimensions with no stride, followed by batch normalization and a LeakyReLu () activation function. The transition output is fed into the decoder, which consists of 4 2D deconvolution blocks. Each comprises a 2D deconvolution layer, batch normalization, and LeakyReLu activation function (). Each output of a deconvolution block is concatenated with the output (concatenated 1D/2D) of the symmetric convolution block. Finally, in the last 2D deconvolution block of the decoder, we use the Tanh activation function to have a signal between and as in the original normalized input signal.
Unlike SOTA approaches, our method directly tackles the challenges of segment recovery and lead reconstruction. To achieve this, it: (1) employs a U-Net architecture that effectively captures horizontal information; (2) offers a simpler optimization process compared to GAN-based approaches; and (3) utilizes a hybrid 1D/2D U-Net architecture to model both inter-lead and intra-lead relationships.
We trained the model for 100 epochs with a batch size of 256. We also used the Adam optimizer with a learning rate of 0.01. The hyperparameter in Equation 1 was tuned by choosing values within the range . The final value of used for the experiments was fixed to (cf. Section A.1).
@@ 3.2. Discussion
We now present the main experimental results to demonstrate the effectiveness of ECGrecover in addressing lead reconstruction and segment recovery. Specifically, we first evaluate the performance of our approach using the metrics outlined in Section 3.1.3. Next, we analyze its ability to preserve key ECG features and assess the reconstructed signal’s impact on inter-lead correlation. Finally, we showcase a real-world application of ECGrecover by applying it also to downstream AI tasks. Further results are provided in Appendix A.
Throughout this section, we use the C notation to indicate the input sample that, at testing time, has been masked according to the considered configuration . Therefore, we compare the reconstructed lead (e.g., lead I from the ECG in output when the input sample was masked with C1) with its original counterpart.
Figure 4 shows examples of lead II reconstructions. The plots suggest that ECGrecover can also correct basal drift, the effect where a signal’s base axis (x-axis) appears to ‘wander’ or move up and down rather than be straight. We elaborate on this in Section A.3.
@@
Lead | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | ECGrecover | EKGAN (joo2023twelve) | Pix2Pix (isola2017image) | CopyPaste | |
@@
Figure 5 presents an evaluation of ECGrecover, alongside benchmark state-of-the-art methods, focusing on PCC, RMSE and DTW for the Generepol dataset. We relegate to Figure 14 the complete sets of the results across all the 17 scenarios and all the 12 leads.
@@
As for the ECG segment-recovery problem, in Figure 5, we specifically focused on configuration C1 as it represents the scenario with primers of shorter duration (i.e., 0.8sec) where most of the signal needs to be reconstructed. Additionally, we analyze the settings C3 and C5 as they reflect real-life situations when digitizing paper ECGs. Regarding the ECG lead-reconstruction problem, we highlight in the figure the configurations C as it is the one considered also considered in (joo2023twelve), and C because of its relevance in wearable device recordings and paper-stored ECGs, where Lead II, akin to Lead I, is predominantly captured. For each of the input masks, we tested the ability of the model to reconstruct the limb leads I, II, and III and the chest leads V1 and V6. Practitioners note that leads II and V1 are useful leads for measuring heart rate.
@@
Overall, ECGrecover outperforms or reaches the same results as the SOTA, regardless of the chosen metric. As depicted in Figure 5, lead V6 emerges as the most straightforward to reconstruct, both in ECG segment-recovery (C1, C3, C5) and ECG lead-reconstruction (CI, CII). Conversely, lead III is the most challenging to reconstruct from the other leads. In particular, while our method consistently delivers satisfactory results across all samples tested, competing methods exhibit more significant variability (e.g., DTW values for EKGAN). Interestingly, the simple CopyPaste method yields reasonable results regarding RMSE and DTW. However, as seen from the PCC results, the method fails to preserve the linear relationships between the original and reconstructed leads, underscoring the method’s limitations as the heart rhythm may vary through time, thus losing sync. This highlights the importance of evaluating the methods from multiple perspectives.
Referring to the comprehensive results in Figure 14, our analysis shows that, in contrast to ECGrecover, the competing approaches EKGAN and Pix2Pix face significantly more challenges in the ECG segment-recovery task compared to the ECG lead-reconstruction task. On the other hand, the CopyPaste method struggles primarily with the lead-reconstruction problem. This outcome is not surprising as EKGAN and Pix2Pix were specifically designed for ECG lead-reconstruction. In contrast, CopyPaste, which essentially duplicates the signal across leads, performs poorly when the lead to reconstruct is less correlated with the primer lead in the mask. We emphasize that in Section A.6, we conduct an additional evaluation of our proposed method using the PTB-XL dataset (wagner2020ptb).
@@ 3.2.2. On the QT distance preservation
An ECG visually represents the heart’s electrical activity during a cardiac cycle, encompassing a single heartbeat, which involves the heart filling with blood and then contracting to pump it out. The standard ECG trace includes several key components, each corresponding to specific electrical events within the heart. The intervals and segments between these components (QT distance, PR interval, etc.) also provide important information about the cardiac cycle’s timing and electrical conduction through the heart. Understanding and evaluating them helps diagnose a wide range of cardiac abnormalities and conditions (parmet2003electrocardiograms). In Table 2, we evaluated the methods’ capacity to generate a signal that can be effectively analyzed. The performance is computed as the difference between the QT distance calculated in the reconstructed signals and the QT distance computed in the original signals (-QT). Specifically, we segmented each ECG into heart-beats by identifying the position of the R-peak within the original leads and isolated the specific spike by taking a 1-second window centered on the R-peak. For detecting the peaks, we use the Neurokit library (Makowski2021neurokit), which supports extracting critical features from the ECG. At each heartbeat, we identified the Q wave and the end of the T wave, and calculate the average QT interval for each ECG.
The results in Table 2 are grouped w.r.t. Sot+ and Sot- true label in the Generepol study. Indeed, as described in Section 3.1.1, in this study participants were administered Sotalol, a drug known to prolong the QT interval and in some cases alter the morphology of the T-wave. These effects make QT interval measurement more challenging (salem2017genome; prifti2021deep). Remarkably, our analysis did not reveal a significant performance difference in ECGrecover between ECGs recorded before and after drug intake. The method demonstrate consistent performance across various input masks and lead reconstructions. While ECGrecover outperforms or reaches the same results as its competitors, the CopyPaste method occasionally achieves better results. The explanation lies in the inherent nature of the CopyPaste technique, which maintains beat information through direct replication. By duplicating each beat identically, the CopyPaste method inherently preserves the average QT distance with lower variability, reflecting an artifact.
@@
We are interested in inter-lead correlation during ECG reconstruction. Specifically, in Figure 6, the matrix entry at index represents PCC between the -th reconstructed lead and the -th original lead, given that the input mask applied is . Thus, the -th row illustrates the model’s effectiveness in utilizing the specific -th lead’s data to reconstruct the remaining leads of the ECG accurately. Conversely, the -th column indicates which input mask most significantly guides the model towards a more accurate reconstruction.
@@
Figure 6 highlights the strong correlation between leads V3, V4, V5 and V6. Lead III exhibits a lower correlation with the other leads. While surprising at first, given that lead III can theoretically be derived from leads I and II through Einthoven’s triangle (parmet2003electrocardiograms), this characteristic has also been observed in (jain2023redundancy) on a distinct dataset. Moreover, similar conclusions can also be found in (zhang202212) regarding lead V1. Their findings revealed a notably low correlation between lead V1 and the other leads, with PCC close to 0, suggesting that the information provided by lead V1 is substantially distinct from that of the other leads. Consequently, it is understandable that the mask C in Figure 6 is the most challenging to reconstruct. Noteworthy, PCC measures linear correlation, and other complex relationships will not be captured. Factors like noise or artifacts in individual physiology can also contribute to lower correlation values.
3.2.4. Reconstructing digitized paper-stored ECGs
PTB-XL (2203 ECGs) | LUDB (200 ECGs) | ||||||
Lead | PCC () | RMSE () | DTW () | PCC () | RMSE () | DTW () | |
C | I | 0.96±0.18 | 0.04±0.04 | 0.0±0.0 | 0.96±0.20 | 0.03±0.02 | 0.0±0.0 |
II | 0.95±0.18 | 0.05±0.06 | 0.0±0.0 | 0.95±0.21 | 0.04±0.04 | 0.0±0.0 | |
III | 0.92±0.21 | 0.05±0.06 | 0.0±0.0 | 0.91±0.22 | 0.04±0.03 | 0.0±0.0 | |
AVL | 0.91±0.23 | 0.05±0.09 | 0.0±0.0 | 0.91±0.24 | 0.06±0.07 | 0.0±0.0 | |
AVR | 0.95±0.21 | 0.05±0.06 | 0.0±0.0 | 0.95±0.22 | 0.04±0.04 | 0.0±0.0 | |
AVF | 0.93±0.21 | 0.05±0.08 | 0.0±0.0 | 0.92±0.21 | 0.05±0.05 | 0.0±0.0 | |
V1 | 0.86±0.31 | 0.09±0.23 | 0.0±0.0 | 0.86±0.31 | 0.07±0.09 | 0.0±0.0 | |
V2 | 0.90±0.32 | 0.17±0.27 | 0.0±0.0 | 0.87±0.36 | 0.08±0.09 | 0.0±0.0 | |
V3 | 0.87±0.31 | 0.27±0.35 | 0.0±0.0 | 0.87±0.30 | 0.13±0.11 | 0.0±0.0 | |
V4 | 0.90±0.29 | 0.21±0.32 | 0.0±0.0 | 0.88±0.31 | 0.09±0.11 | 0.0±0.0 | |
V5 | 0.93±0.28 | 0.13±0.23 | 0.0±0.0 | 0.92±0.32 | 0.07±0.09 | 0.0±0.0 | |
V6 | 0.94±0.28 | 0.09±0.28 | 0.0±0.0 | 0.94±0.30 | 0.06±0.07 | 0.0±0.0 |
Historically, ECGs have been printed and stored on paper or saved as images (Einthoven1903), which presents clear storage and exploitation challenges, as highlighted in Section 1. In the most common formats for paper-stored ECGs, each lead is only partially represented, typically showing only 2.5 or 5 seconds of the usual 10-second signal. This section evaluates the proposed method in a real-life scenario by examining ECGrecover’s ability to reconstruct the missing parts of ECGs that were initially stored on paper and digitized.
We considered the LUDB (kalyakulina2020ludb) and PTB-XL (wagner2020ptb) datasets for the experiments. Specifically, these datasets are available in both the original paper-stored version (in .pdf) and the corresponding numerical version (in .xml). The ECGs saved in .pdf format was digitized with ECGminer software (santamonica2024ecgminer) to produce a partial .xml version. The .pdf version of the ECGs corresponds to configuration C3, except for lead II, which is complete. We refer to this new setting as C. In contrast, the .xml version includes the full 12-lead tracks, making it well-suited for comprehensive testing of the reconstruction method. We evaluated ECGrecover, trained on the Generepol dataset, to assess its robustness in reconstructing ECGs from various databases. To align with the training phase, the signals were preprocessed in the same way as presented in Section 3.1.2.
The results presented in Table 3 demonstrate that the proposed method achieves strong performance in this setting. Specifically, for the PTB-XL database, the method yields an overall mean PCC of 0.92, an RMSE of 0.1, and an SDTW of 0. For the LUDB database, it achieves an average PCC of 0.91, an RMSE of 0.06, and a DTW distance of 0. These positive outcomes can be attributed to the carefully designed configuration we sought to reconstruct. In particular, Creal-life incorporates the full lead II along with partial versions of other leads. This configuration serves as an intermediate setup between C3 and C, providing the model with a richer and more diverse set of input information compared to the training configurations. The enhanced input likely accounts for the improved predictive performance. Notably, as illustrated in Figure 6, lead II alone demonstrates the ability to reconstruct all ECG leads with an accuracy of 0.87. By integrating additional information from other leads, the hybrid configuration in C likely amplifies the model’s performance, resulting in the observed improvements.
3.2.5. Applying ECGrecover to AI-based applications
Here we illustrate the use of ECGrecover in a downstream AI-based application task, predicting the risk of Torsade-de-Pointes events. As detailed in Section 3.1.1, Sot- and Sot+ denote ECGs from healthy individuals before and after ingesting 80mg of Sotalol. Thus, we utilize the CNN model initially crafted by Prifti et al. (prifti2021deep) for TdP risk prediction. This DenseNet model comprises six blocks, each containing eight dense convolutional layers. We trained it with the Adam optimizer, employing a learning rate of 0.001 and a dropout rate of 0.2. We evaluate our method as a data augmentation technique.
During training, for a given input sample , we augmented the dataset with , recovered from configuration C, with a 0.25 probability and trained the TdP-risk model for 50 epochs, iterating the training phase five times with seeds [0-4] for the random number generator. We focused on configurations C3 and C5, commonly encountered in real-life settings, as discussed in Section 3.2.1. The TdP-risk model’s accuracy on the original testing set without data augmentation is 92.47%. In contrast, with data augmentation, the accuracy is 93.84%±0.0065 for C3, 93.73%±0.002 for C5, and 92.17%±0.0128 for C. Results are presented as mean +/- standard deviation across the five training stages. The results show that data regenerated with ECGrecover can be used to augement the initial dataset, thus improving model performance. In particular, data regenerated from C3 increases model performance by 1%.